OpenStudy (anonymous):

During a class experiment, students were required to dilute 50.0 mL of 2.50 M Na2SO4 by mixing it with 50.0 mL of distilled water. One of the groups accidentally mixed the 2.50 M Na2SO4 solution with 50.0 mL of 0.0500 M Na2SO4 instead of distilled water. What is the molarity of the resulting solution that this group made? A. 1.25 M B. 1.28 M C. 2.45 M D. 2.55 M

OpenStudy (australopithecus):

Molarity = Moles/Litres

OpenStudy (australopithecus):

So what you need to do is figure out total moles, find moles of Na2SO4 in 50mL=0.05L of 2.50M solution find moles of Na2SO4 in 50mL=0.05L of 0.05M solution Then simply add the moles from both solutions and then add the volume then calculate the new molarity

OpenStudy (australopithecus):

Using the formula I provided you with

OpenStudy (anonymous):

I'm trying to get the answer ... not work it out

OpenStudy (australopithecus):

Oh well go to yahoo answers, I'm not here to do your homework for you.

OpenStudy (australopithecus):

or you could save time and just work it out yourself because it is an easy question.

OpenStudy (anonymous):

its an easy question? maybe for you but not for me. such an ignorant person you are. bye

OpenStudy (australopithecus):

You undersell yourself, I'm 100% certain you could solve this problem

OpenStudy (australopithecus):

Molarity = Moles/Litres Molarity*Litres = Moles Just plug values into these equations and you have your answer

OpenStudy (australopithecus):

You should not be so quick to insult someone who is trying to help you.

OpenStudy (australopithecus):

You could even show me your work and I dont mind tell you if you are on the right track.

OpenStudy (mathmate):

If you are mixing two solutions of different concentrations, you need to calculate the total number of moles (C1V1+C2V2) and divide by the total volume (V1+V2). Example, Mix 30.0 mL of 0.200M H2SO4 with 25.0 mL of 0.100M H2SO4. Find final molarity. #mol in 30 mL of 0.2M = 0.03 L *0.2 mol/L= 0.006 mol #mol in 25 mL of 0.1M = 0.025 L * 0.1 mol/L = 0.0025 mol Total #mol of H2SO4 = 0.005+0.0025 = 0.0075 mol Total volume = 0.030+0.025 = 0.055 L Final molarity = (use @Australopihecus 's formula) 0.0075 mol / 0.055 L = 0.136 mol/L = 0.136M REMEMBER: This is an example to help you, this is NOT the answer. Go ahead to work on your problem so that you can solve a similar problem later on.